A large retailer obtains merchandise under the credit terms of 2/15, net 30, but routinely takes 50 days to pay its bills. (Because the retailer is an important customer, suppliers allow the firm to stretch its credit terms.) What is the retailer's effective cost of trade credit? Assume a 365-day year. Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Cost of trade credit: =(1+D=(1-D))(365:n)-1 Dis discount rate n is Number of days after discount period Effective cost of tra

*Please rate thumbs up